### v1

# A Flat Minima Perspective on Understanding Augmentations and Model Robustness

Abstract
Model robustness indicates a model’s capability to generalize well on unforeseen distributional shifts, including data corruption, adversarial attacks, and domain shifts. Data augmentation is one of the prevalent and effective ways to enhance robustness. Despite the great success of augmentations in different fields, a general theoretical understanding of their efficacy in improving model robustness is lacking. We offer a unified theoretical framework to clarify how augmentations can enhance model robustness through the lens of loss surface flatness and PAC generalization bound. Our work diverges from prior studies in that our analysis i) broadly encompasses much of the existing augmentation methods, and ii) is not limited to specific types of distribution shifts like adversarial attacks. We confirm our theories through simulations on the existing common corruption and adversarial robustness benchmarks based on the CIFAR and ImageNet datasets, as well as domain generalization benchmarks including PACS and OfficeHome.
1 Introduction
Model robustness, which is a critical factor of deep models in applications requiring high reliability such as autonomous vehicles and medical diagnosis, entails maintaining performance against data distribution shifts. In the past decade, data augmentation has been widely used as a popular and pragmatic technique to enhance the model performance, as well as the robustness against data corruption, adversarial attacks, or even domain shifts augment_survey ; ME-ADA ; Mixup:2018 ; DeepAug10:2021 ; AutoAugment . The intuition of its efficacy relies on the belief that augmentations enrich the training data distribution, which allows models to easily extrapolate to unseen data distributions, which is the so-called generalization capability.
Figure 1: A conceptual overview of how augmentations relate to model robustness. Theorem 1 shows that the neighboring region of an input datapoint can be mapped to an equivalent region in the parameter space (and vice versa). Theorem 2 demonstrates that ensuring nonzero probability for an augmentation within this neighborhood can promote flatter minima in parameter space by effectively regulating the proximal regions of both the input and parameter spaces. Finally, Theorem 3 presents a generalization bound for models trained with augmentations, asserting that augmentations with dense representations near the original image foster a tighter generalization gap.
Despite the utility of augmentations, a unified theory formally clarifying how augmentations enhance model robustness against diverse distribution shifts has not been well established. Prior analyses are mostly limited to either specific augmentations or adversarial robustness ME-ADA ; ADA ; AugmentNAdvGood ; NoLabelNAdvGood1 ; NoLabelNAdvGood2 ; aug_analysis_adv1 . Although some studies step forward from the aforementioned works, their analyses rely solely on empirical observations without providing a theoretical rationale, or they focus exclusively on certain categories of distribution shifts aug_analysis_cc1 ; aug_analysis_custom1 ; aug_analysis_dg . For instance, zhang2024duality links adversarial training with sharpness-aware minimization, but the results are confined to adversarial robustness.
In this paper, we first present a set of theoretical explanations revealing how data augmentations can reinforce model robustness across general distributional changes in the perspective of flat minima. Our analysis has two main branches: i) First, we formally bridge the general form of augmentations to the improved generalization bound. To give a brief sketch of our development, we start to demonstrate the equivalence between the input space region covered by the augmented samples and the corresponding parameter space region with the same loss values (formalized by Theorem 1). Based on the equivalence, we then claim that the augmentations flatten the loss surface on the parameter space (Theorem 2), finally reaching to the improved generalization bound against distribution shifts via leveraging the flattened loss surface (Theorem 3). ii) Next, we evaluate existing augmentation methods across a range of robustness benchmarks, encompassing data distribution shifts caused by common corruptions, adversarial attacks, and domain changes. Our findings show that augmentations with sample coverage near the original image—which aligns with our assumption (Assumption 1)—tend to consistently enhance overall model robustness. In contrast, augmentations that rarely improve robustness under diverging distribution shifts exhibit vacant density in close proximity to the original image.
2 Preliminaries
2.1 Basic Notations
Let us consider an input $x\in\mathbb{R}^{n}$ from input space $\mathcal{X}$, which is paired with a target label $y\in\mathbb{R}^{c}$ from label space $\mathcal{Y}$, where $n$ and $c$ are the dimensions of the input space and the label space. A model $f(\cdot;\theta):\mathbb{R}^{n}\rightarrow\mathbb{R}^{c}$ parameterized by $\theta\in\mathbb{R}^{p}$ maps a given input to the estimated label, where $p$ is the dimension of the model parameter space. The loss function $\mathcal{L}(\cdot,\cdot)$ quantifies how far the predicted label $f(x;\theta)$ deviates from the true label $y$, expressed as $\mathcal{L}\big{(}f(x;\theta),y\big{)}$. Given a dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$ of $N$ samples drawn from data distribution $\mathcal{D}$, the true risk and the empirical risk are:
$\mathcal{E}_{\mathcal{D}}(\theta):=\mathbb{E}_{(x,y)\sim\mathcal{D}}\big{[}\mathcal{L}(f(x;\theta),y)\big{]},\quad\hat{\mathcal{E}}_{\mathcal{D}}(\theta):=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f(x_{i};\theta),y_{i})$ (1)
respectively, where $\mathcal{E}_{\mathcal{D}}(\theta)$ and ^$\hat{\mathcal{E}}_{\mathcal{D}}(\theta)$ denote the true and the empirical risk.
2.1.1 Data Augmentation
Data augmentation $\mathcal{A}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ augments a given input $x$ to augmented input $\tilde{x}:=\mathcal{A}(x)$. Let us further represent it with the difference between the original input and the augmented one: $\tilde{x}:=\mathcal{A}(x)=x+\delta$, where $\delta\in\mathbb{R}^{n}$.  [^1] By formulating augmentation as an additive perturbation, we transform the perturbation on the parameter space to the one on the input space. Details are in the following section. Also, we define $\mathcal{P}_{\mathcal{A}}(\tilde{x}|x)$ as the probability density function of an augmented sample $\tilde{x}$ given $x$. Finally, we let $\tilde{\mathcal{D}}:=\{(\mathcal{A}(x),y):(x,y)\sim\mathcal{D}\}$ represent the augmented dataset.
2.2 Model Flatness
2.2.1 Definition
Model flatness characterizes the extent of change in the model’s loss values across proximate points in the parameter space. When the loss rapidly changes around the found minima, it indicates that the model is located at sharp minima. Otherwise, it denotes flat minima when the loss varies smoothly. The change of losses around the model parameters can be formalized as follows:
$\max_{\|\Delta\|\leq\gamma}\mathbb{E}_{(x,y)\sim\mathcal{D}}\big{[}\mathcal{L}(f(x;\theta+\Delta),y)-\mathcal{L}(f(x;\theta),y)\big{]},$ (2)
where $\Delta\in\mathbb{R}^{p}$ perturbs model parameter $\theta$ within radius $\gamma>0$ to maximize the loss.
2.2.2 Sharpness-aware Minimization
The most popular principal way for finding flat minima is utilizing sharpness-aware minimization (SAM) optimizer SAM:2021 , which formally transforms loss minimization into a min-max optimization:
$\min_{\theta}\max_{\|\Delta\|\leq\gamma}\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}(f(x;\theta+\Delta),y)],$ (3)
As formulated in the work of SAM:2021 , when adding the loss value at $\theta$, only the first term of (2) remains, yielding the objective function in (3). Through the minimization of the maximization term, SAM aims to find the minima $\theta$ with flatter loss surfaces around $\gamma$ radius.
3 How Can Augmentations Enhance Robustness?
In this section, we provide a rigorous link between data augmentations and improved generalization capability. Our main claims are twofold: i) Translation between equivalent input manifold and parameter manifold, ii) Association of flatter loss in the input space via augmentations with flatness in the parameter space, leading to a reduced generalization gap.
3.1 Translation Between Input and Parameter Perturbations
Consider a perturbation $\delta\in\mathbb{R}^{n}$ around an input vector $x$. Intuitively, there should be a corresponding perturbation $\Delta\in\mathbb{R}^{p}$ around the parameter $\theta$ that yields the same output, i.e., $f(x+\delta;\theta)=f(x;\theta+\Delta)$. Moreover, for any perturbation in the closed ball $\{\delta\in\mathbb{R}^{n}:\|\delta\|\leq\gamma\}$ in the input space, we hypothesize an equivalent or compensatory region $\mathcal{C}_{\Theta}$ in the parameter space—that is, for every $\|\delta\|\leq\gamma$, there is a matching parameter perturbation $\Delta\in\mathcal{C}_{\Theta}$ producing the same output.
Conversely, the same intuition in reverse applies for parameter perturbations. Consider a perturbation $\Delta\in\mathbb{R}^{p}$ around $\theta$. For any $\|\Delta\|\leq\gamma$, we posit an analogous compensatory region $\mathcal{C}_{\mathcal{X}}$ in the input space such that $f(x;\theta+\Delta)=f(x+\delta;\theta)$. Specifically, there is some input shift $\delta\in\mathcal{C}_{\mathcal{X}}$ that replicates the same output change induced by $\Delta$.
Our goal is to formalize this duality of perturbations in both the input and parameter spaces. We now present the notion of functional compensatory sets, which identifies the precise sets of input shifts needed to compensate for every parameter perturbation of norm up to $\gamma$, and vise versa.
3.1.1 Functional Compensatory Sets
Definition 1.
(Functional Compensatory Sets)
(A) Parameter-to-Input. Given a dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ and a parameter perturbation set $\{\Delta\in\mathbb{R}^{p}:\|\Delta\|\leq\gamma\},$ a functional compensatory set $\mathcal{C}_{\mathcal{X}}$ in the input space satisfies:
[…]
$f(x;\theta+\Delta)=f(x+\delta;\theta).$
This set is not necessarily unique, and it characterizes all possible input shifts that replicate the effect of parameter perturbations within norm $\gamma$.
(B) Input-to-Parameter. Given a dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ and an input perturbation set $\{\delta\in\mathbb{R}^{n}:\|\delta\|\leq\gamma\},$ a functional compensatory set $\mathcal{C}_{\Theta}$ in the parameter space satisfies:
[…]
$f(x+\delta;\theta)=f(x;\theta+\Delta).$
This set $\mathcal{C}_{\Theta}$ need not be unique, and captures all possible parameter shifts that replicate the effect of any input perturbation within norm $\gamma$.
3.1.2 Formalizing the Input–Parameter Perturbation Duality
The following result shows that an $n$-dimensional ball in the input space, whose radius depends on dataset $\mathcal{D}$ and the Jacobian’s singular values, can compensate for any parameter perturbation $\|\Delta\|\leq\gamma$. Conversely, the same analysis guarantees that a $p$-dimensional ball in the parameter space, determined by $\mathcal{D}$ and the Jacobian, can compensate for any input perturbation $\|\delta\|\leq\ \gamma$.
Let $\mathbf{J}$ be the Jacobian of a function $f:\mathbb{R}^{n}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{c}$, where $\mathbb{R}^{n},\mathbb{R}^{p},$ and $\mathbb{R}^{c}$ are the input, parameter, and output spaces. The Jacobian matrix $\mathbf{J}\in\mathbb{R}^{c\times(n+p)}$ can be decomposed into its input side and the parameter side partial derivatives, i.e. $\mathbf{J}=\left[\mathbf{J_{x}}\ \ \mathbf{J}_{\boldsymbol{\theta}}\right]$, where $\mathbf{J_{x}}\in\mathbb{R}^{c\times n}$ and $\mathbf{J}_{\boldsymbol{\theta}}\in\mathbb{R}^{c\times p}$. We will use the notation $\mathbf{J_{x}}(x_{i})$ and $\mathbf{J}_{\boldsymbol{\theta}}(\theta)$ when we need to explicitly represent the evaluation of $\mathbf{J_{x}}$ and $\mathbf{J}_{\boldsymbol{\theta}}$ at point $x_{i}\in\mathbb{R}^{n}$ and $\theta\in\mathbb{R}^{p}$, respectively. Finally, let $\sigma_{\min}^{x_{i}}$ and $\sigma_{\max}^{\theta}$ represent the mininum and the maximum values of the matrix $\mathbf{J_{x}}(x_{i})$ and $\mathbf{J}_{\boldsymbol{\theta}}(\theta)$. Leveraging Sard’s theorem Sard together with Taylor expansion, we construct two sets—$\mathcal{S}_{\mathcal{X}}$ and $\mathcal{S}_{\Theta}$—that fulfill the definition of functional compensatory sets $\mathcal{C}_{\mathcal{X}}$ and $\mathcal{C}_{\Theta}$ as follows:
Theorem 1.
(Input–Parameter Duality via Functional Compensatory Sets) Consider a function $f:\mathbb{R}^{n}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{c}$.
(A) Parameter-to-Input Define the ball of parameter perturbations $\{\Delta\in\mathbb{R}^{p}:\|\Delta\|\leq\gamma\}$ and
$\mathcal{S}_{\mathcal{X}}\;=\;\left\{\delta\in\mathbb{R}^{n}:\|\delta\|\;\leq\;\max_{x\in\mathcal{D}}\!\Bigl{(}\dfrac{\sigma_{\max}^{\theta}}{\sigma_{\min}^{x}}\Bigr{)}\,\gamma\right\}.$
Then, $\mathcal{S}_{\mathcal{X}}$ is always a functional compensatory set for the parameter perturbation ball and dataset $\mathcal{D}$.
(B) Input-to-Parameter Conversely, define the ball of input perturbations $\{\delta\in\mathbb{R}^{n}:\|\delta\|\leq\gamma\}$ and
$\mathcal{S}_{\Theta}\;=\;\left\{\Delta\in\mathbb{R}^{p}:\|\Delta\|\;\leq\;\max_{x\in\mathcal{D}}\!\Bigl{(}\dfrac{\sigma_{\max}^{x}}{\sigma_{\min}^{\theta}}\Bigr{)}\,\gamma\right\}.$
Then, $\mathcal{S}_{\Theta}$ is always a functional compensatory set for the input perturbation ball and dataset $\mathcal{D}$.
Proof.
See Appendix B.1. ∎
Remark 1.1.
(Loss Stability Across Input and Parameter Spaces) Ensuring the model’s loss is low on all datapoints within a small neighborhood around $x$ (i.e., for all $x+\delta$ with $\delta\in\mathcal{S}_{\mathcal{X}}$) guarantees that the corresponding neighborhood in parameter space $\{\theta+\Delta:\|\Delta\|\leq\gamma\}$ also has low loss. Conversely, if the model’s loss is low within a small neighborhood around the parameters (i.e. $\theta+\Delta$ for $\Delta\in\mathcal{S}_{\Theta}$), then any corresponding input perturbation $\delta$ with $\|\delta\|\leq\gamma$ also yields low loss. Hence, stability in the input space induces stability in the parameter space, and vice versa.
3.2 Linking Augmentations to Model Flatness and Generalization
Built upon the theorems above, we now formalize how can augmentations lead to flatter loss landscape around the minima and the improved generalization capability. Before providing the details, let us rephrase the formal definition of flat minima, which is called $b$-flat minima shi2021overcoming :
Definition 2.
($b$-flat local minima) Given loss $\mathcal{L}(\cdot,\cdot)$ and dataset $\mathcal{D}$, a model parameter $\theta\in\mathbb{R}^{p}$ is $b$-flat minima if the followings hold for the perturbation on parameter $\Delta\in\mathbb{R}^{p}$:
$\text{For all}\>\|\Delta\|\leq b,\>\>\hat{\mathcal{E}}_{\mathcal{D}}(\theta+\Delta)=\hat{\mathcal{E}}_{\mathcal{D}}(\theta)\text{ and }$
$\text{For some}\>\|\Delta\|>b,\>\>\hat{\mathcal{E}}_{\mathcal{D}}(\theta+\Delta)>\hat{\mathcal{E}}_{\mathcal{D}}(\theta).$
For input $x$, an augmentation $\mathcal{A}$ induces a distribution of variants $\tilde{x}$ near $x$. Let $\gamma_{\mathcal{A}}$ be the largest radius of a ball around $x$ with non-zero probability. If $\gamma_{\mathcal{A}}>0$, we say $\mathcal{A}$ covers the local neighborhood. We assume this property for augmentations currently considered and show in Section 4 that it is critical for the robustness of augmentations.
Assumption 1.
Given $x\in\mathbb{R}^{n}$ and augmentation $\mathcal{A}(\cdot)$, the probability density function $\mathcal{P}_{\mathcal{A}}(\tilde{x}|x)$ satisfies:
$\text{For all}\>\|\delta\|\leq\gamma_{\mathcal{A}},\>\mathcal{P}_{\mathcal{A}}(\tilde{x}|x)>0,$ (4)
where $\delta=\tilde{x}-x$ and $\gamma_{\mathcal{A}}$ is some positive real number.
Let $\tilde{\mathcal{D}}$ represent the augmented dataset with respect to $\mathcal{D}$ and $\gamma_{\mathcal{A}}$ the radius value satisfying Assumption 1. Now, let $\Theta^{*}$ and $\tilde{\Theta}^{*}$ be the sets of the optimal model parameters whose elements $\theta^{*}\in\Theta^{*}$ and $\tilde{\theta}^{*}\in\tilde{\Theta}^{*}$ satisfy the following equalities, $\mathcal{E}_{\mathcal{D}}(\theta^{*})=0$ and $\mathcal{E}_{\mathcal{\tilde{D}}}(\tilde{\theta}^{*})=0$, respectively. Then our claim is that the minimum $b$-flatness among the solution parameters in $\tilde{\Theta}^{*}$, shows large $b$ (flatter) than the minimum $b$-flatness among the solutions in $\Theta^{*}$, which is trained on $\mathcal{D}$:
Theorem 2.
(Flatness of $\tilde{\theta}^{*}$) Let $\theta^{*}\in\Theta^{*}$ and $\tilde{\theta}^{*}\in\tilde{\Theta}^{*}$ be $b^{*}$ and $\tilde{b}^{*}$-flat minima, respectively. The following inequality holds:
$\min_{\theta^{*}\in\Theta^{*}}b^{*}\leq\min_{\tilde{\theta}^{*}\in\tilde{\Theta}^{*}}\tilde{b}^{*}.$
Proof.
See Appendix B.2. ∎
Remark 2.1.
(Any augmentation with nearby representations around the original promotes flatter minima) The key understanding of the theorem above is that if augmentations cover the ball-shaped region around the given original sample with radius $\gamma_{\mathcal{A}}$, then the model $\tilde{\theta}^{*}$ suppresses the loss values of the region. Next, the flat-region on the input space is translated to the functional compensatory set region on parameter space, which at least contains $\{\delta:\|\delta\|\leq\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}\}$.
The final linkage to the generalization capability is straightforward by relying on the prior theoretical results that bridge Robust Risk Minimization (RRM) and the generalization bound SWAD:2021 . RRM aims to find a flat minimum, where its surrounding region shows flat loss surface. Based on Theorem 2, augmentations make the optimal parameter locate at a flatter surface, and it implies the improved generalization bound (referred to Theorem 3). Let us define $\gamma_{\Theta}:=\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}$, then the following theorem holds:
Theorem 3.
(Generalization bound) Given $M$ covering sets $\{\Theta_{k}\}_{k=1}^{M}$ of parameter space $\Theta$ with $\Theta=\bigcup_{k=1}^{M}\Theta_{k}$ and $\text{diam}(\Theta)=\sup_{\theta,\theta^{\prime}\in\Theta}||\theta-\theta^{\prime}||_{2}$, where $M=\left\lceil\frac{\text{diam}(\Theta)}{\gamma_{\Theta}}\right\rceil^{d}$, and VC dimension $v_{k}$ for each $\Theta_{k}$, the following inequality holds with probability at least $1-\delta$:
$\mathcal{E}_{\mathcal{T}}(\tilde{\theta}^{*})$ <^$<\hat{\mathcal{E}}_{\mathcal{\tilde{D}}}(\tilde{\theta}^{*})+\frac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\max_{k\in[1,M]}\left[\sqrt{\frac{v_{k}\log(N/v_{k})}{N}+\frac{\log(M/\delta)}{N}}\right],$
where $\textbf{Div}(\mathcal{D},\mathcal{T})=2\sup_{A}|\mathbb{P}_{\mathcal{D}}(A)-\mathbb{P}_{\mathcal{T}}(A)|$ measures the maximal discrepancy between the source and the target distributions $\mathcal{D}$ and $\mathcal{T}$, and $N$ is the number of samples drawn from $\mathcal{\tilde{D}}$.
Proof.
See Appendix B.3. ∎
Remark 3.1.
(Augmentations improve generalization against data distribution shifts) The theorem implies that the minimization of empirical loss for the augmented dataset, i.e., ^$\hat{\mathcal{E}}_{\mathcal{\tilde{D}}}(\tilde{\theta}^{*})$, directly aims the tighter generalization bound on target distribution $\mathcal{T}$. Also, the term $M$ is related to $\gamma_{\mathcal{A}}$, which measures how augmented data distribution covers a wider range (referred to Assumption 1). When augmentation covers a wider range around the original sample, i.e., a larger $\gamma_{\mathcal{A}}$, it suppresses $M$, leading to the smaller last term of generalization bound. These interpretation elucidates how augmentations enhance the generalization capability against unseen data distribution shifts.
4 Empirical Examination on Proximal Density, Model Flatness, and Generalization
In this section, we put our theory to test by examining how augmentations influence model robustness. Earlier, we established that any augmentation assigning non-zero probability to samples lying in a neighborhood around the original image (Assumption 1) maps to an analogous region in parameter space (Theorem 1), which in turn induces flatter minima (Theorem 2) and yields a tighter generalization bound (Theorem 3). Therefore, augmentations with rich proximal density should yield strong robustness gains under diverse distribution shifts, whereas augmentations that fail to enhance robustness should suffer from near-zero representations in the proximal region of the original image.
4.1 Augmentation Methods to Be Considered
Following the taxonomy of augment_survey , augmentations can be categorized into model-free, model-based, and policy-based augmentations. Among the immense existing augmentation methods, we focus on well-known or advanced methods rather than basic or elementary augmentations. For the model-free augmentation type, we select AugMix AugMix9:2021 and PixMix PixMix:2022 , which are advanced forms of mixing multiple clean images to the original image. For the model-based augmentation type, we choose StyleAug StyleAug and DeepAug DeepAug10:2021 , which utilize image-to-image models to diversify the style of the clean images. Finally, among the policy-based augmentation type, we consider RandAugment RandAugment , which learns the policy for augmenting images.
4.2 Tests on Proximal Density
We first measure each augmentation’s proximity (or “proximal density”) to the original image via observing the empirical cumulative distribution function (eCDF) of $L_{2}$ distances. Specifically, we compute how quickly the eCDF accumulates at smaller values, which provides a clear indicator of each augmentation’s density near the original image. Formally, if $\{d_{i}\}_{i=1}^{N}$ are these distances, then the eCDF is defined as $F_{N}(\gamma)=\tfrac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{d_{i}\leq\gamma\}$. High values in the lower-distance region of $F_{N}(\gamma)$ signifies that a large fraction of augmented samples remain close to the original images. Conversely, an augmentation whose representations are spread out will exhibit small values to small gamma, indicating sparser coverage.
As illustrated in Figure 2 and Table 1, on CIFAR-100 dataset AugMix have dense representations near the original, RandAug and PixMix show sparse but non-zero clustering, and DeepAug and StyleAug exhibit nearly zero coverage at low distances. On the ImageNet dataset, the results are analogous to CIFAR, except that DeepAug’s eCDF now accumulates much more rapidly and has dense representations in the proximal region unlike CIFAR.
Figure 2: Distance histograms on CIFAR-100 illustrating augmentation density (top: entire distribution, bottom: near origin). From the highest to the lowest coverage near the origin: AugMix is dense, RandAug and PixMix are sparse, and DeepAug and StyleAug have zero. A dense concentration of histogram close to the original suggests a higher likelihood that $\gamma_{\mathcal{A}}\gg 0$, whereas an absence of samples in the low-distance region indicates $\gamma_{\mathcal{A}}=0$ (referring to Assumption 1).
Table 1: Empirical CDF $\big{(}F_{N}(\gamma)\big{)}$ for selected distance thresholds $(\gamma)$, evaluated across various augmentation methods on CIFAR-10/100 and ImageNet. Red entries indicate zero or negligible probability $(<5\cdot 10^{-5})$, and green entries exceed 0.01, illustrating how densely each augmentation stays near the original image.
Distance CIFAR-100 Distance ImageNet AugMix RandAug PixMix DeepAug StyleAug AugMix RandAug PixMix DeepAug StyleAug $\gamma=0.01$ 0.0024 0.0060 0.0011 0.0000 0.0000 $\gamma=20.0$ 0.0005 0.0001 0.0002 0.0005 0.0000 $\gamma=0.05$ 0.0107 0.0060 0.0011 0.0000 0.0000 $\gamma=30.0$ 0.0031 0.0009 0.0013 0.0028 0.0000 $\gamma=0.10$ 0.0202 0.0061 0.0012 0.0000 0.0000 $\gamma=40.0$ 0.0113 0.0034 0.0054 0.0110 0.0001 $\gamma=0.50$ 0.0993 0.0378 0.0012 0.0001 0.0000 $\gamma=50.0$ 0.0317 0.0104 0.0169 0.0311 0.0008
4.3 Tests on Model Flatness
We now explore how an augmentation’s density near the original image influences flatness in the learned model parameters, linking our results from Theorems 1 and 2 with empirical flatness.
Flatness metrics. We assess each model using three quantitative measures: PAC-Bayesian flatness measure $\mu_{\text{PAC-bayes}}$ PACBayes , local sharpness $\epsilon_{\text{sharp}}$ eps-sharpness , and Loss-Pass Filter (LPF) metric, a strong indicator of generalization LPF . Each of these metrics reflects the difference between the “proximal loss” around a model parameter and its original loss, which aligns well with our theoretical framework. We therefore avoid purely “pinpoint-based” measures (e.g., maximum Hessian eigenvalues or their traces), as they do not fully capture our theorems’ emphasis on local neighborhoods in parameter space. For more comprehensive descriptions of each metric, refer to Appendix C.1.1.
Experimental results. We train WideResNet-40-2 architecture on CIFAR-10/100 datasets adopting the five augmentations. For further details on training, see Appendix C.1.1. Table 2 presents each model’s flatness on CIFAR, under different augmentation settings. In line with our theorems and assumption, methods that retain nonzero probability of producing samples close to the original image (AugMix, RandAug, PixMix) yield flatter minima than ERM, irrespective of whether their coverage is dense or sparse. Meanwhile, augmentations that fail to exceed ERM in achieving flatter minima (DeepAug, StyleAug) are those exhibiting near-zero proximal density. These observations align with Theorems 1, 2, and Assumption 1, underscoring that maintaining a nontrivial concentration of augmented samples near the original input is a critical aspect for finding flatter minima.
In the (tiny)ImageNet setting, augmentations with high proximal density (Table 1) again produce substantially flatter minima than plain ERM. All methods except StyleAug achieve this benefit, and DeepAug does so as well in this case. Detailed table statistics are at Appendix C.1.
Table 2: Flatness metrics of ERM and different augmentations on CIFAR benchmarks (:absent:$\downarrow$:↓ : The lower the better, i.e., flatter minimum). Overall, augmentations with nonzero probability near the original image consistently yield flatter minima, whereas those with near-zero coverage struggle to outperform ERM on all the flatness measures.
Dataset Metrics      ERM      AugMix      RandAug      PixMix      DeepAug      StyleAug CIFAR-10 $\mu_{\text{PAC-Bayes}}\downarrow$ 166.72 116.70 (-50.02) 113.68 (-53.04) 105.30 (-61.42) 66.89 (-99.83) 122.76 (-43.96) LPF $\downarrow$ 1.01 0.52 (-0.49) 0.46 (-0.55) 0.37 (-0.64) 0.71 (-0.30) 1.14 (+0.13) $\epsilon_{\text{sharp}}$downarrow 10.28 7.42 (-2.86) 7.32 (-2.96) 9.19 (-1.09) 11.13 (+0.85) 19.84 (+9.56) CIFAR-100 $\mu_{\text{PAC-Bayes}}\downarrow$ 179.92 155.74 (-24.18) 151.15 (-28.77) 136.07 (-43.85) 71.56 (-108.36) 135.48 (-44.44) LPF $\downarrow$ 2.37 1.87 (-0.50) 1.67 (-0.70) 1.40 (-0.97) 2.38 (+0.01) 3.39 (+1.02) $\epsilon_{\text{sharp}}$downarrow 8.98 8.35 (-0.63) 7.95 (-1.03) 8.00 (-0.98) 13.49 (+4.51) 42.82 (+33.84)
4.4 Tests on Generalization Capability
We finally investigate whether augmentations having rich proximal density indeed fosters model robustness against diverse distribution shifts (Theorem 3).
Benchmarks to be considered. Generalization refers to a model’s ability to perform well on unseen data, which here spans natural corruptions of clean samples, adversarial attacks, and distributional changes from domain shifts. Given that prior work has already explored simpler train-test generalization scenarios (with limited distributional discrepancies), we focus on more challenging shifts, such as common corruptions and adversarial attacks on CIFAR CIFAR , tinyImageNet, and ImageNet IN-C7:2019 , as well as domain generalization using PACS PACS and OfficeHome OfficeHome .
Robustness Benchmarks and Metrics. Following the setup in Table 2, we train our models on CIFAR-10/100, tinyImageNet, and ImageNet datasets and then evaluate their robustness to common corruptions and untargeted PGD $L_{2}/L_{\infty}$ adversarial attacks. For domain generalization, we adopt the approach of DomainBed DomainBed , employing ResNet50 and reporting average accuracies over three independent runs for each benchmark and augmentation strategy. In Table 3, lower values represent better performance on corruption and adversarial benchmarks, while higher values indicate better domain generalization. See Appendix C.1.2 for further details on benchmarks, metrics, and training.
Experimental Results. The outcomes align closely with Theorems 1 to 3. Specifically, every augmentation method that maintained nonzero probability near the original image surpasses ERM in overall robustness. Moreover, the more densely each method populate the proximity region, the larger the performance boost. For instance, AugMix populates many augmented samples near the originals in both CIFAR and tinyImageNet, which translates into top-level robustness. RandAug and PixMix keep only a few proximal variants, enough to outperform ERM but not to dominate. DeepAug leaves proximal region empty in CIFAR, yet fills the region succesfully on ImageNet. Therfore, DeepAug underperforms on CIFAR while excelling on (tiny)ImageNet. StyleAug produces no nearby samples at all and yields no robustness gains.
We interpret these observations as the evidence of the relationship between augmentations and robustness strongly holding, when the nonzero probability condition (Assumption 1) is met.The results also co-aligns with our theorems, highlighting the importance of rich proximal density in improving model generalization capability.
Table 3: Comparison of various augmentation methods against ERM across multiple robustness scenarios, including common corruptions (CIFAR-10/100-C/$\overline{\text{C}}$, tinyImageNet-C, ImageNet-C), adversarial attacks ($L_{2}$, $L_{\infty}$), and domain shifts (PACS, OfficeHome). Metrics with ()($\downarrow$)( ↓ ) favor lower values (corruptions and adversarial) and ()($\uparrow$)( ↑ ) favor higher values (domain generalization). Differences from ERM are shown in parentheses.
Robustness Benchmark ERM      AugMix      RandAug      PixMix      DeepAug      StyleAug CIFAR-10-C $\downarrow$ 30.54 15.24 (-15.30) 19.65 (-10.89) 10.57 (-19.97) 31.92 (+1.38) 30.50 (-0.04) CIFAR-10-$\overline{\text{C}}$downarrow 8.13 1.73 (-6.40) 3.07 (-5.06) 1.05 (-7.08) 17.99 (+9.86) 9.34 (+1.21) Common CIFAR-100-C $\downarrow$ 59.05 42.64 (-16.41) 46.59 (-12.46) 35.14 (-23.91) 62.34 (+3.29) 70.92 (+11.87) Corruption CIFAR-100-$\overline{\text{C}}$downarrow 18.32 5.01 (-13.31) 8.31 (-10.01) 4.78 (-13.54) 39.54 (+21.22) 30.77 (+12.45) tinyImageNet-C $\downarrow$ 74.58 63.47 (-11.11) 66.95 (-7.63) 61.51 (-13.07) 62.51 (-12.07) 79.31 (+4.73) ImageNet-C $\downarrow$ 69.37 69.34 (-0.03) 66.88 (-2.49) 61.51 (-7.86) 56.68 (-12.69) 75.52 (+6.15) CIFAR-10, $L_{2}$downarrow 77.67 70.68 (-6.99) 76.30 (-1.37) 65.24 (-12.43) 91.07 (+13.40) 91.13 (+13.46) CIFAR-10, $L_{\infty}$downarrow 98.48 99.57 (+1.09) 99.86 (+1.38) 99.46 (+0.98) 100.00 (+1.52) 99.97 (+1.49) CIFAR-100, $L_{2}$downarrow 98.72 92.78 (-5.94) 95.92 (-2.60) 90.11 (-8.61) 98.37 (-0.35) 99.67 (+0.95) Adversarial CIFAR-100, $L_{\infty}$downarrow 99.94 99.67 (-0.27) 99.93 (-0.01) 99.60 (-0.34) 99.99 (+0.05) 99.98 (+0.04) Attacks tinyImageNet $L_{2}$downarrow 57.89 61.12 (+3.23) 58.86 (+0.97) 63.57 (+5.68) 52.91 (-4.98) 83.50 (+25.61) tinyImageNet $L_{\infty}$downarrow 99.91 99.89 (-0.02) 99.95 (+0.04) 99.94 (+0.03) 99.92 (+0.01) 100.00 (+0.09) ImageNet, $L_{2}$downarrow 76.16 76.06 (-0.10) 76.09 (-0.07) 75.44 (-0.72) 75.37 (-0.79) 87.71 (+11.55) ImageNet, $L_{\infty}$downarrow 99.70 99.69 (-0.01) 99.70 (+0.00) 99.63 (-0.07) 99.62 (-0.08) 99.92 (+0.22) Domain PACS $\uparrow$ 79.8 82.1 (+2.3) 83.4 (+3.6) 79.5 (-0.3) 64.4 (-15.4) 83.7 (+3.9) Shifts OfficeHome $\uparrow$ 68.2 68.5 (+0.3) 67.4 (-0.8) 63.7 (-4.5) 44.7 (-23.5) 64.5 (-3.7)
5 Discussion and Future Work
In this section, we outline the strengths, limitations, and potential directions for future research stemming from our work. The primary contributions of our study are:
  • We first offer a flat minima perspective to elucidate how augmentations affect robustness.
    • While prior research has explored the relationship between augmentations and robustness, to the best of our knowledge, no existing work has utilized the concept of flat minima to establish this connection.
    • A weak duality between adversarial training (AT) and sharpness-aware minimization (SAM) is known to exist zhang2024duality . Nonetheless, the result holds only for linear models with Gaussian data and fails to cover standard augmentations.
  • We first establish a formal connection between the general form of data augmentations and enhanced robustness under general distribution shifts.
    • Previous studies faced limitations such as focusing solely on: i) experimental analysis, ii) specific augmentations, or iii) certain distributional shifts.
    • In contrast, our analysis imposes no such restrictions on augmentation types or benchmarks. We validate our findings through experiments across diverse robustness benchmarks.
Despite these contributions, our work has certain limitations. Specifically, we have not conducted massive experiments to validate our theorem across a broader range of benchmarks and augmentation methods. For instance:
  • Our theoretical framework does not encompass label-manipulating augmentations like Mixup.
  • Sophisticated adversarial attacks such as the C&W attack remain unexplored in our experiments.
Future research may advance our work by incorporating label-manipulating augmentations and examining a more extensive set of techniques. For further discussion of theorem applications, flat-minima optimizers, and adversarial training, see Appendix E, which also provides interpretation to augmentation compositions based on our findings.
6 Related Works
In this section, we provide a brief overview of previous works on data augmentation and model robustness. For more detailed discussions on prior works, refer to Appendix D.
Augmentation Strategies. Augmentations can be grouped as model-free, model-based, or policy-based augmentations augment_survey . Model-free methods act on one image (translation, rotation, masking CutOut ; HideNSeek ) or several images (Mixup, CutMix Mixup:2018 ; CutMix:2019 ). Advanced augmentations (AugMix, PixMix AugMix9:2021 ; PixMix:2022 ) consisting of multiple image augmentation steps, are known to improve corruption robustness AugMix9:2021 ; PixMix:2022 . Model-based approaches rely on pretrained generators (CGAN and its variants CGAN ; CGAN_var1 ; CGAN_var2 ; CGAN_var3 ; CGAN_var4 ) or image-to-image networks (DeepAugment, ANT, StyleAug) to generate augmented images. Policy-based schemes learn augmentation policies: AutoAugment AutoAugment and its successors FastAA ; FasterAA ; RandAugment ; AdaTransform ; AdvAA ; AugMax fall into this group. Despite such augmentations’ strong empirical gains, a unifying theory on understanding their impacts on robustness still missing.
Augmentation and Robustness. A number of previous studies tries to find the link between augmentation and robustness in several angles. A game-theoretic view exists on interpreting augmentation’s effect on common corruption robustness aug_analysis_cc1 . A Fourier perspective on defining adversarial and common corruption robustness exists AugFourier . Interpretations on adversarial perturbation as the worst-case distribution shifts in the data ME-ADA ; ADA exist. Certain augmentations with weight averaging is known to improve adversarial and common corruption robustness AugmentNAdvGood ; boundary_thickness . Yet, these explanations either remain empirical or tied to specific augmentations, lacking applicability in general settings.
Beyond Augmentations. The duality between adversarial training and sharpness-aware minimization optimizer is known to exist on a linear classifier over Gaussian distributions zhang2024duality . Preserving fairness under distribution shifts by mimicking the shift as perturbations utilizing transportation function has been studied jiang2023chasing . Flatter minima is shown to mitigate boundary tilting problem boundary_tilting and is empirically found to be strong against common corruptions flat_minima_cc . Still, none of these works can explain augmentation-driven robustness.
7 Conclusion
Our framework links data augmentation to robustness through the lens of flat minima under general distribution shifts. When augmentations densely fill the neighborhood of an image, they map to broad, flat regions in parameter space, tightening generalization bounds and improving robustness to distribution shifts. Sparse augmentations, by contrast, leave gaps near the original image and confer little protection. Experiments on CIFAR, ImageNet, PACS, and OfficeHome—covering common corruptions, adversarial attacks, and domain shifts—corroborate our theory. We hope these insights guide the design of next-generation augmentation methods, including those built on foundation models.
8 Acknowledgment
This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II201336, Artificial Intelligence Graduate School Program (UNIST)), (No. IITP-2025-RS-2022-00156361, Innovative Human Resource Development for Local Intellectualization program), and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00459023).
Appendix A Supplementary Materials
This appendix contains additional material that could not be incorporated into the main paper due to page constraints, including detailed proofs of the theorems and experimental details. Section B provides detailed proofs of Theorems 1 to 3, Section C presents experimental details and additional results, Section D addresses related works in detail, and Section E offers further discussion on broader applicability of our findings including practical usage.
Appendix B Proofs on Theorems
B.1 Proof on Theorem 1
Figure 3: An illustration of Theorem 1 (A). Given a parameter perturbation set $\{\Delta\in\mathbb{R}^{p}:\|\Delta\|\leq\gamma\}$ and datapoints $x_{1},...,x_{n}$, each $x_{i}$ has a minimal functional compensatory region $\mathcal{S}_{min}(x_{i})$ shaped like an ellipsoid. All of the ellipsoids can be collectively enclosed by an isotropic ball $\mathcal{S}_{\mathcal{X}}$, whose radius is determined by the largest principal axis across the ellipsoids. The exact reverse process is done for the Theorem 1 (B).
B.1.1 Proof on Theorem 1.A
Theorem 1 (A). (Parameter-to-Input) Consider a function $f:\mathbb{R}^{n}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{c}$. For any $\gamma>0$, define the ball of parameter perturbations $\{\Delta\in\mathbb{R}^{p}:\|\Delta\|\leq\gamma\}$ and
$\mathcal{S}_{\mathcal{X}}\;=\;\left\{\delta\in\mathbb{R}^{n}:\|\delta\|\;\leq\;\max_{x\in\mathcal{D}}\!\Bigl{(}\dfrac{\sigma_{\max}^{\theta}}{\sigma_{\min}^{x}}\Bigr{)}\,\gamma\right\}.$
Then, $\mathcal{S}_{\mathcal{X}}$ is always a functional compensatory set for the parameter perturbation ball and dataset $\mathcal{D}$.
Proof.
As a brief recap regarding notations on Jacobian matrices, $\mathbf{J}$ is the Jacobian of a function $f:\mathbb{R}^{n}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{c}$, and the Jacobian matrix $\mathbf{J}$ can be expressed as its input and parameter side differentials, i.e. $\mathbf{J}=\left[\mathbf{J_{x}}\ \ \mathbf{J}_{\boldsymbol{\theta}}\right]$. The notation $\mathbf{J_{x}}(x_{i})$ and $\mathbf{J}_{\boldsymbol{\theta}}(\theta)$ will be occasionally used to represent the evaluation of $\mathbf{J_{x}}$ and $\mathbf{J}_{\boldsymbol{\theta}}$ at point $x_{i}\in\mathbb{R}^{n}$ and $\theta\in\mathbb{R}^{p}$.
Using the first-order Taylor expansion, $f(x;\theta+\Delta)\approx f(x;\theta)+\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta$. This approximation is supported by two main observations. First, when $f$ is a linear model, the Taylor expansion becomes exact, and a linear model with infinite width serves as a universal approximator. Second, supported by lazy training regime [8, 13, 35, 3] and Neural Tangent Kernel (NTK) [25, 32] statements, training loss for a deep model rapidly converges to zero with minimal parameter updates. The neural network is known to behave like its linearization around the initial weight in the entire training process, and the change in Jacobian becomes negligible.
In the same vein, $f(x+\delta;\theta)\approx f(x;\theta)+\mathbf{J_{x}}\cdot\delta$. Given a dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ and $\gamma>0$, our interest is finding a functional compensatory set $\mathcal{C}_{\mathcal{X}}$ such that for any $x\in\mathcal{D}$ and $\|\Delta\|\leq\gamma$, there exists $\delta\in\mathcal{C}_{\mathcal{X}}$ satisfying $f(x;\theta+\Delta)=f(x+\delta;\theta)$. Utilizing the above two Taylor expansions, the problem reduces to finding the region of $\delta$ satisfying the equation
$\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta=\mathbf{J_{x}}\cdot\delta$ (6)
By Sard’s theorem [45]—which states that the set of critical values of a smooth function has Lebesgue measure zero—the Jacobian matrices have full rank almost everywhere. Therefore, we will assume the matrices $\mathbf{J}_{\mathbf{x}}$ and $\mathbf{J}_{\boldsymbol{\theta}}$ to be full-rank matrices as in [58, 14]. Since $c\ll n,p$ and $\mathbf{J}_{\boldsymbol{\theta}},\mathbf{J_{x}}$ are full-rank, there always exists a solution $\delta$ for the equation (6). Using psuedoinverse notation $\mathbf{J}_{\mathbf{x}}^{+}$ and kernel notation $\text{ker}(\mathbf{J_{x}})$ for $\mathbf{J_{x}}$, the solution for equation (6) can be written as
$\delta=\mathbf{J}_{\mathbf{x}}^{+}\cdot\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta+w$
, where $w\in\text{ker}(\mathbf{J_{x}})$. Under the constraint $\|\Delta\|\leq\gamma$, the full feasible solution region can be written as $\mathcal{S}_{\max}(x):=\{\mathbf{J}_{\mathbf{x}}^{+}\mathbf{J}_{\boldsymbol{\theta}}\,\Delta+w:\|\Delta\|\leq\gamma,w\in\text{ker}(\mathbf{J_{x}})\}$, and the minimal covering solution region as $\mathcal{S}_{\min}(x):=\{\mathbf{J}_{\mathbf{x}}^{+}\mathbf{J}_{\boldsymbol{\theta}}\,\Delta:\|\Delta\|\leq\gamma\}$. By the definition of functional compensatory set, both $\mathcal{S}_{\max}(x)$ and $\mathcal{S}_{\min}(x)$ can be the functional compensatory sets for $x\in\mathcal{D}$ and $\|\Delta\|\leq\gamma$. For the sake of simplicity, we will investigate the feasible region $\mathcal{S}_{\min}(x)$.
There are two things to consider at this point. First, the set $\{Ax:\|x\|\leq\gamma\}$ is always a rotated ellipsoid, whether the matrix $A$ is degenerate or not. Thus, $\mathcal{S}_{\min}(x):=\{\mathbf{J}_{\mathbf{x}}^{+}\mathbf{J}_{\boldsymbol{\theta}}\,\Delta:\|\Delta\|\leq\gamma\}$ is an $c$-dimensional rotated ellipsoid having singular values of $\mathbf{J}_{\mathbf{x}}^{+}\mathbf{J}_{\boldsymbol{\theta}}$ as the lengths of its principal semi-axes. Second, any set $\mathcal{S}$ containing $\mathcal{S}_{\min}$ is also a compensatory set since $\mathcal{S}_{\min}\subseteq\mathcal{S}$ implies that for all $\delta\in\mathcal{S}_{\min}$, there exists $\delta\in\mathcal{S}$. We will derive the closed-form region $\mathcal{S}$ using the fact that given an ellipsoid with maximum possible principal semi-axes length $l_{\max}$, the ellipsoid in encapsulated by an isotropic ball with radius $l_{\max}$ as well.
Let $\mathbf{J_{x}}=U_{x}\Sigma_{x}V_{x}^{T}$ and $\mathbf{J}_{\boldsymbol{\theta}}=U_{\theta}\Sigma_{\theta}V_{\theta}^{T}$ be the singular value decomposition representations, and $\mathbf{J}_{\mathbf{x}}^{+}=V_{x}\Sigma_{x}^{+}U_{x}^{T}$ be the psuedoinverse representation of $\mathbf{J_{x}}$. Let $\sigma_{1}^{x},...,\sigma_{c}^{x}$ and $\sigma_{1}^{\theta},...,\sigma_{c}^{\theta}$ be the singular values of $\mathbf{J_{x}}$ and $\mathbf{J}_{\boldsymbol{\theta}}$ (i.e. diagonal elements of $\Sigma_{x}$ and $\Sigma_{\theta}$.) Trivially, $\mathbf{J_{x}}^{+}$ has singular values $1/\sigma_{1}^{x},...,1/\sigma_{c}^{x}$.
$\|\mathbf{J}_{\mathbf{x}}^{+}\mathbf{J}_{\boldsymbol{\theta}}\,\Delta\|$ $\leq\|\mathbf{J}_{\mathbf{x}}^{+}\|\cdot\|\mathbf{J}_{\boldsymbol{\theta}}\|\cdot\|\Delta\|$ $(\because\|AB\|\leq\|A\|\cdot\|B\|)$
$=\|V_{x}\Sigma_{x}^{+}U_{x}^{T}\|\cdot\|U_{\theta}\Sigma_{\theta}V_{\theta}^{T}\|\cdot\gamma$
$\leq\|\Sigma_{x}^{+}\|\|\Sigma_{\theta}\|\cdot\gamma$ $(\because\|Q\|=1\text{ for orthonormal matrix }Q)$
$=\dfrac{\max_{i=1,...,c}\ \sigma_{i}^{\theta}}{\min_{i=1,...,c}\ \sigma_{i}^{x}}\cdot\gamma=:\dfrac{\sigma_{\max}^{\theta}}{\sigma_{\min}^{x}}\cdot\gamma$
Since the above inequality is about arbitrary datapoint $x\in\mathcal{D}$, we can write inequalities for the whole dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ as
$\|\mathbf{J}_{\mathbf{x}}(x_{1})^{+}\mathbf{J}_{\boldsymbol{\theta}}(\theta)\,\Delta\|\leq\sigma_{\max}^{\theta}/\sigma_{\min}^{x_{1}}\cdot\gamma,\cdots,\|\mathbf{J}_{\mathbf{x}}(x_{N})^{+}\mathbf{J}_{\boldsymbol{\theta}}(\theta)\,\Delta\|\leq\sigma_{\max}^{\theta}/\sigma_{\min}^{x_{N}}\cdot\gamma.$
Then, the following bound holds:
$\text{For all }x\in\mathcal{D}\text{ and }\|\Delta\|\leq\gamma,\quad\|\mathbf{J}_{\mathbf{x}}^{+}(x)\mathbf{J}_{\boldsymbol{\theta}}(\theta)\,\Delta\|\leq\max_{x\in\mathcal{D}}\dfrac{\sigma_{\max}^{\theta}}{\sigma_{\min}^{x}}\cdot\gamma.$
Therefore, $\mathcal{S}_{\mathcal{X}}\;=\;\left\{\,\delta\in\mathbb{R}^{n}:\|\delta\|\;\leq\;\max_{x\in\mathcal{D}}\left(\dfrac{\sigma_{\max}^{\theta}}{\sigma_{\min}^{x}}\right)\cdot\gamma\right\}$ is a functional compensatory set for parameter perturbations $\{\Delta\in\mathbb{R}^{p}:\|\Delta\|\leq\gamma\}$.
B.1.2 Proof on Theorem 1.B
Theorem 1 (B). (Input-to-Parameter) Consider a function $f:\mathbb{R}^{n}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{c}$. Conversely, for any $\gamma>0$, define the ball of input perturbations $\{\delta\in\mathbb{R}^{n}:\|\delta\|\leq\gamma\}$ and
$\mathcal{S}_{\Theta}\;=\;\left\{\Delta\in\mathbb{R}^{p}:\|\Delta\|\;\leq\;\max_{x\in\mathcal{D}}\!\Bigl{(}\dfrac{\sigma_{\max}^{x}}{\sigma_{\min}^{\theta}}\Bigr{)}\,\gamma\right\}.$
Then, $\mathcal{S}_{\Theta}$ is always a functional compensatory set for the input perturbation ball and dataset $\mathcal{D}$.
Proof.
We employ the same strategy to that in the proof of Theorem 1.A. (Appendix B.1.1) For conciseness, we do not restate the complete set of assumptions and justifications presented previously. Under the first-order Talyor expansion, $f(x;\theta+\Delta)\approx f(x;\theta)+\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta$ and $f(x+\delta;\theta)\approx f(x;\theta)+\mathbf{J_{x}}\cdot\delta$, resulting in $\mathbf{J}_{\boldsymbol{\theta}}\cdot\Delta=\mathbf{J_{x}}\cdot\delta$. Since $\mathbf{J}_{\boldsymbol{\theta}}$ and $\mathbf{J_{x}}$ are full-rank, the solution for the equation becomes
$\Delta=\mathbf{J}_{\boldsymbol{\theta}}^{+}\cdot\mathbf{J_{x}}\cdot\delta+w$
, where $w\in\text{ker}(\mathbf{J}_{\boldsymbol{\theta}})$. Under the constraint $\|\delta\|\leq\gamma$, the minimal feasible solution region becomes $\mathcal{S}_{\min}(x):=\{\mathbf{J}_{\boldsymbol{\theta}}^{+}\mathbf{J_{x}}\,\delta:\|\delta\|\leq\gamma\}$. Under the singular value decomposition, $\mathbf{J}_{\boldsymbol{\theta}}=U_{\theta}\Sigma_{\theta}V_{\theta}^{T}$ and $\mathbf{J_{x}}=U_{x}\Sigma_{x}V_{x}^{T}$. Let $\sigma_{1}^{\theta},...,\sigma_{c}^{\theta}$ and $\sigma_{1}^{x},...,\sigma_{c}^{x}$ represent the singular values of $\mathbf{J}_{\boldsymbol{\theta}}$ and $\mathbf{J_{x}}$, respectively. Then,
$\|\mathbf{J}_{\boldsymbol{\theta}}^{+}\mathbf{J_{x}}\,\delta\|$ $\leq\|\mathbf{J}_{\boldsymbol{\theta}}^{+}\|\cdot\|\mathbf{J_{x}}\|\cdot\|\delta\|$ $(\because\|AB\|\leq\|A\|\cdot\|B\|)$
$=\|V_{\theta}\Sigma_{\theta}^{+}U_{\theta}^{T}\|\cdot\|U_{x}\Sigma_{x}V_{x}^{T}\|\cdot\gamma$
$\leq\|\Sigma_{\theta}^{+}\|\|\Sigma_{x}\|\cdot\gamma$ $(\because\|Q\|=1\text{ for orthonormal matrix }Q)$
$=\dfrac{\max_{i=1,...,c}\ \sigma_{i}^{x}}{\min_{i=1,...,c}\ \sigma_{i}^{\theta}}\cdot\gamma=:\dfrac{\sigma_{\max}^{x}}{\sigma_{\min}^{\theta}}\cdot\gamma$
Since this holds for any $x\in\mathcal{D}$, we can write for $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$:
$\|\mathbf{J}_{\mathbf{\theta}}(\theta)^{+}\mathbf{J_{x}}(x_{1})\,\delta\|\leq\sigma_{\max}^{x_{1}}/\sigma_{\min}^{\theta}\cdot\gamma,\cdots,\|\mathbf{J}_{\mathbf{\theta}}(\theta)^{+}\mathbf{J}_{\mathbf{x}}(x_{N})\,\delta\|\leq\sigma_{\max}^{x_{N}}/\sigma_{\min}^{\theta}\cdot\gamma.$
Leading to:
$\text{For all }x\in\mathcal{D}\text{ and }\|\delta\|\leq\gamma,\quad\|\mathbf{J}_{\mathbf{x}}^{+}(x)\mathbf{J}_{\boldsymbol{\theta}}(\theta)\,\delta\|\leq\max_{x\in\mathcal{D}}\dfrac{\sigma_{\max}^{x}}{\sigma_{\min}^{\theta}}\cdot\gamma.$
Consequently, $\mathcal{S}_{\Theta}\;=\;\left\{\,\Delta\in\mathbb{R}^{p}:\|\Delta\|\;\leq\;\max_{x\in\mathcal{D}}\left(\dfrac{\sigma_{\max}^{x}}{\sigma_{\min}^{\theta}}\right)\cdot\gamma\right\}$ is a functional compensatory set for input perturbations $\{\delta\in\mathbb{R}^{n}:\|\delta\|\leq\gamma\}$.
B.2 Proof on Theorem 2
Theorem 2.
(Flatness of $\tilde{\theta}^{*}$) Let $\theta^{*}\in\Theta^{*}$ and $\tilde{\theta}^{*}\in\tilde{\Theta}^{*}$ be $b^{*}$ and $\tilde{b}^{*}$-flat minima, respectively. The following inequality holds:
$\min_{\theta^{*}\in\Theta^{*}}b^{*}\leq\min_{\tilde{\theta}^{*}\in\tilde{\Theta}^{*}}\tilde{b}^{*}.$ (9)
Proof.
We consider $f(\cdot;\theta)$ to be a universal approximator for functions from the input space $\mathbb{R}^{n}$ to the output space $\mathbb{R}^{c}$. For the sake of simplicity, we will omit the subscript notations $\theta^{*}\in\Theta^{*}$ and $\tilde{\theta}^{*}\in\tilde{\Theta}^{*}$ in $\min_{\theta^{*}\in\Theta^{*}}$ and $\min_{\tilde{\theta}^{*}\in\tilde{\Theta}^{*}}$ henceforth. We prove $\min b^{*}=0<\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}\leq\min\tilde{b}^{*}.$
Let $\theta^{*}_{0}$ be an optimal parameter of a model that satisfies for all $(x,y)\in\mathcal{D},$ $\mathcal{L}(f(x;\theta_{0}^{*}),y)$ $=0$ and for all $(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{c}\ /\ \mathcal{D},$ $\mathcal{L}(f(x;\theta_{0}^{*}),y)>0$. Let $b_{0}^{*}$ denote the $b$-value of the $b$-flat minima for $\theta^{*}_{0}$.
Suppose $b_{0}^{*}>0$. By the definition of $b$-flat minima, for all $\|\Delta\|\leq b_{0}^{*}$ and $(x,y)\in\mathcal{D}$, $\mathcal{L}(f(x;\tilde{\theta}^{*}_{0}+\Delta),y)=\mathcal{L}(f(x;\tilde{\theta}^{*}_{0}),y))$. Define $\mathcal{C}_{\Theta}:=\{\Delta\in\mathbb{R}^{p}:\|\Delta\|\leq b_{0}^{*}\}$. By Theorem 1, $\mathcal{C}_{\Theta}$ is a functional compensatory set for input perturbations
$\left\{\delta\in\mathbb{R}^{n}:\left(\max_{x\in\mathcal{D}}\dfrac{\sigma_{\max}^{x}}{\sigma_{\min}^{\theta}}\right)^{-1}\cdot b_{0}^{*}\right\}=:\left\{\delta\in\mathbb{R}^{n}:\lambda\cdot b_{0}^{*}\right\}$
[…]
This implies $\min b^{*}\leq\min b_{0}^{*}\leq 0$, i.e. $\min b^{*}=0.$
We now show $\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}\leq\min\tilde{b}^{*}$ as follows:
Let $\tilde{\mathcal{D}}=\{(\mathcal{A}(x),y):(x,y)\in\mathcal{D}\}$ represent an augmented dataset. Let $\tilde{\theta}^{*}$ be some optimal parameter which achieves zero loss on the augmented dataset, i.e. for all $(\tilde{x},y)\in\tilde{\mathcal{D}},$ $\mathcal{L}(f(\tilde{x};\tilde{\theta}^{*}),y))=0$. Finally, let $\tilde{b}^{*}$ be the $b$ value for the optimal parameter of $\tilde{\mathcal{D}}$. We show
$\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}\big{/}\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}\leq\min\tilde{b}^{*}$
as follows. By definition, $\gamma_{\mathcal{A}}$ represents the maximum $L_{2}$ distance that augmentation has non-zero probabilities around the original input. Define $\mathcal{C}_{\mathcal{X}}=\{\delta\in\mathbb{R}^{n}:\|\delta\|\leq\gamma_{\mathcal{A}}\}.$ Under Theorem 1, $\mathcal{C}_{\mathcal{X}}$ is the compensatory set for the parameter perturbations
$\left\{\Delta\in\mathbb{R}^{p}:\max_{x\in\mathcal{D}}\left(\dfrac{\sigma_{\max}^{x}}{\sigma_{\min}^{\theta}}\right)^{-1}\gamma_{\mathcal{A}}\right\}.$
By definition of $b$-flat minima, we have $\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}\leq\min\tilde{b}^{*}$. ∎
B.3 Proof on Theorem 3
Theorem 3.
(Generalization bound) Given $M$ covering sets $\{\Theta_{k}\}_{k=1}^{M}$ of parameter space $\Theta$ with $\Theta=\bigcup_{k=1}^{M}\Theta_{k}$ and $\text{diam}(\Theta)=\sup_{\theta,\theta^{\prime}\in\Theta}||\theta-\theta^{\prime}||_{2}$, where $M=\left\lceil\frac{\text{diam}(\Theta)}{\gamma_{\Theta}}\right\rceil^{d}$, and VC dimension $v_{k}$ for each $\Theta_{k}$, the following inequality holds with probability at least $1-\delta$:
$\mathcal{E}_{\mathcal{T}}(\tilde{\theta}^{*})$ <^$<\hat{\mathcal{E}}_{\mathcal{\tilde{D}}}(\tilde{\theta}^{*})+\frac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\max_{k\in[1,M]}\left[\sqrt{\frac{v_{k}\log(N/v_{k})}{N}+\frac{\log(M/\delta)}{N}}\right],$
where $\textbf{Div}(\mathcal{D},\mathcal{T})=2\sup_{A}|\mathbb{P}_{\mathcal{D}}(A)-\mathbb{P}_{\mathcal{T}}(A)|$ measures the maximal discrepancy between the source and the target distributions $\mathcal{D}$ and $\mathcal{T}$, and $N$ is the number of samples drawn from $\mathcal{\tilde{D}}$.
Proof.
Remind that $\mathcal{\tilde{D}}$ represents augmented data distribution, and $\gamma_{\Theta}:=\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}$ the radius of the compensatory set $\mathcal{C}_{\Theta}$ defined in Theorem 1. The robust empirical risk is originally defined [7] as:
^$\hat{\mathcal{E}}_{\mathcal{D}}^{\gamma}(\theta):=\max_{\|\delta\|\leq\gamma}\frac{1}{N}\sum_{i=1}^{N}\big{[}L(f(x_{i},\theta+\delta),y_{i})\big{]}.$
Analogously, we define the counterpart in the input space as follows:
^$\hat{\mathscr{E}}_{\mathcal{D}}^{\gamma}(\theta):=\max_{\|\delta\|\leq\gamma}\frac{1}{N}\sum_{i=1}^{N}\big{[}L(f(x_{i}+\delta,\theta),y_{i})\big{]}.$
For the sake of readability, the notation $\mathcal{B}(M,N,\delta):=\max_{k\in[1,M]}\left[\sqrt{\frac{v_{k}\log(N/v_{k})}{N}+\frac{\log(M/\delta)}{N}}\right]$ will be used henceforth.
We now list the necessary propositions from [7] for the rest of the proof.
Proposition 1. For all $\theta\in\Theta,\ |\mathcal{E}_{\mathcal{D}}(\theta)-\mathcal{E}_{\mathcal{T}}(\theta)|\leq\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})$
Proposition 2. For all $\theta\in\Theta,\gamma>0,\ \mathcal{E}_{\mathcal{D}}(\theta)\leq\hat{\mathcal{E}}_{\mathcal{D}}^{\gamma}(\theta)+\mathcal{B}(M,N,\theta)$.
We are now ready to prove the rest of our theorem.
$\mathcal{E}_{\mathcal{T}}(\tilde{\theta}^{*})$ $\leq\mathcal{E}_{\mathcal{D}}(\tilde{\theta}^{*})+\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})$ Prep. 1
$\leq\mathcal{E}_{\mathcal{D}}^{\gamma_{\Theta}}(\tilde{\theta}^{*})+\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\mathcal{B}(M,N,\delta)$ Prep. 2, $\gamma=\gamma_{\Theta}$
^$\leq\hat{\mathscr{E}}_{\mathcal{D}}^{\gamma_{\mathcal{A}}}(\tilde{\theta}^{*})+\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\mathcal{B}(M,N,\delta)$ Def. of $\mathscr{E},\gamma_{\Theta},$ Thm. 1
^$\leq\hat{\mathcal{E}}_{\tilde{\mathcal{D}}}(\tilde{\theta}^{*})+\dfrac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\mathcal{B}(M,N,\delta)$ Def. of $\gamma_{\mathcal{A}},\,\tilde{\theta}^{*}$
Appendix C Experimental Details and Additional Experiments
In this section, we provide a comprehensive account of the experiments corresponding to the tables and figures presented in the main paper, along with supplementary results and discussions that were excluded from the primary manuscript due to page constraints. The additional results include extra robustness evaluations on different backbone networks, flatness evaluations on (tiny)ImageNet belonging to Table 2, clean accuracies of models in Table 3, and singular value bounds on Theorem 1.
C.1 Detailed Experimental Methodology
For all the experiments assessing flatness, mean accuracy, mean corruption error (mCE), and adversarial robustness, we trained three independent models per benchmark and reported the averaged results, with the exception of the ImageNet experiment in Table 3, which was conducted only once.
C.1.1 Flatness Metrics
[…]
Training Details. Our training procedure closely follows [21], with the only difference being that we train the WideResNet-40-2 model for 200 epochs instead of 100. We employ an SGD optimizer with an initial learning rate of 0.1, a weight decay of 0.0005, and a momentum of 0.9. Additionally, we use a cosine learning rate decay scheduler, starting with a learning rate of 0.1 and a weight decay of 0.0005. Finally, we evaluate and report the flatness of the converged models.
C.1.2 Robustness Benchmark Details
Domain Generalization Benchmarks. We conducted comprehensive domain generalization experiments on the PACS and OfficeHome datasets to evaluate model robustness and generalizability using augmentations. These datasets are specifically designed to test a model’s ability to generalize across different domains, as they consist of images that share the same labels but differ in their domain representations. For example, the PACS dataset contains images from four distinct domains—’art painting’, ’cartoon’, ’photo’, and ’sketch’—all featuring the same classes such as ’dog’, ’elephant’, and others. In all experiments, models were trained until convergence to ensure fair and consistent comparisons across different settings.
Our baseline employed Empirical Risk Minimization (ERM) with the best hyperparameters selected from DomainBed’s ERM training configuration [17]. To specifically assess the impact of augmentations on model performance, we excluded the heavy augmentation compositions used in the original DomainBed configuration.
We have adopted 2,500 steps of training iterations, which results in slightly faster training than the default setting of 5,000. Specifically for StyleAug, we performed additional tuning of its hyperparameter $\alpha$ alongside the training epochs. Through manual experimentation, we found that setting $\alpha=1.0$ yielded the best results in terms of model robustness and generalization performance.
Common Corruption Benchmarks. We assessed the common corruption robustness of models trained with augmentations on the CIFAR-10/100, tinyImageNet, and ImageNet datasets. Specifically, we utilized their corrupted counterparts—CIFAR-10/100-C/$\overline{\text{C}}$, tinyImageNet-C, and ImageNet-C— which are standard benchmarks generated from the original datasets by applying 15 distinct corruption types at 5 different severity levels. These corruption types include brightness changes, contrast alterations, defocus blur, elastic transformations, fog addition, frost addition, Gaussian blur, glass distortion, impulse noise, JPEG compression, motion blur, pixelation, shot noise, snow addition, and zoom blur. Models were trained and validated on the respective clean datasets, with robustness evaluations performed at test time only to simulate real-world scenarios where models encounter unforeseen corruptions. To quantify model performance across all corruptions and severity levels, we calculated the mean Corruption Error (mCE), which averages the error rates over the different conditions.
For CIFAR-10/100 datasets, the training details are exactly the same as in C.1.1. For tinyImageNet and ImageNet experiments, we have majorly adopted the configurations from [21], with changes in epochs and backbone networks for faster training. We have trained ResNet18 on tinyImageNet and ImageNet for 50 epochs, instead of training ResNet50 for 90 epochs. SGD optimizer with initial learning rate 0.01, weight decay 0.0005, and momentum 0.9 has been adopted. All the models have been trained until convergence, with cosine learning rate decaying scheduling.
Regarding the augmentation techniques, default configuration parameters were used for all baseline augmentations except StyleAug, which had hyperparameter $\alpha$ set to $0.95$ for the CIFAR experiments, and 1.0 for the tinyImageNet experiments.
Adversarial Robustness Benchmarks. To evaluate the robustness of models in adversarial settings, we trained models with each augmentation on CIFAR-10/100, tinyImageNet, and ImageNet, and conducted tests using untargeted Projected Gradient Descent (PGD) attacks based on the $L_{2}$ and $L_{\infty}$ norm. The training environment holds the same as in the common corruption benchmarks. We will use $\epsilon$ and $\alpha$ to denote the maximum allowed perturbation size.
Regarding CIFAR-10/100 experiments, we utilized attack configurations from [55]. Specifically, we deployed PGD-20 $L_{2}$ attacks with $\epsilon=0.5,\alpha=0.0125$ and PGD-7 $L_{\infty}$ attacks with $\epsilon=8/255,\alpha=2/255$. For tinyImageNet experiments, we utilized PGD-10 $L_{2}$ attacks with $\epsilon=0.25,\alpha=0.025$ and PGD-5 $L_{\infty}$ attacks with $\epsilon=8/255,\alpha=2/255$. On ImageNet experiments, we deployed PGD-10 $L_{2}$ attacks with $\epsilon=0.25,\alpha=0.025$ and PGD-2 $L_{\infty}$ attacks with $\epsilon=2/255,\alpha=1/255$. In essence, milder attack configurations have been deployed for tinyImageNet and ImageNet.
C.1.3 Results on Theorem 1’s Bound
Using the CIFAR-100 dataset, we evaluated the singular values for Theorem 1 and obtained $\sigma_{\text{min}}^{x}=0.20,\sigma_{\text{max}}^{x}=13.48,\sigma_{\text{min}}^{\theta}=4.55,\sigma_{\text{max}}^{\theta}=52.56$. The minimum singular values are not excessively small and maximum values not excessively large, thereby preventing loose bounds.
C.1.4 Flatness Evaluations on (tiny)ImageNet
Table 4: Flatness metrics of ERM and different augmentations on CIFAR benchmarks (:absent:$\downarrow$:↓ : The lower the better, i.e., flatter minimum). The results are consistent with Table 1; All the augmentations populating meaningful representations near the original found flatter minima compared to ERM (AugMix, RandAug, PixMix, DeepAug.)
Dataset Metrics      ERM      AugMix      RandAug      PixMix      DeepAug      StyleAug tinyImageNet $\mu_{\text{PAC-Bayes}}\downarrow$ 137.97 113.97 (-24.00) 124.82 (-13.15) 113.97 (-24.00) 137.97 (-0.00) 97.09 (-16.88) LPF $\downarrow$ 1.91 1.10 (-0.81) 1.10 (-0.81) 1.05 (-0.86) 1.53 (-0.38) 1.47 (-0.44) $\epsilon_{\text{sharp}}$downarrow 1.46 0.64 (-0.78) 0.77 (-0.69) 0.55 (-0.91) 0.56 (-0.90) 1.95 (+0.49) ImageNet $\mu_{\text{PAC-Bayes}}\downarrow$ 481.93 481.93 (-0.00) 470.81 (-11.12) 405.52 (-76.41) 405.52 (-76.41) 694.44 (+212.51) LPF $\downarrow$ 0.36 0.35 (-0.01) 0.40 (+0.04) 0.35 (-0.01) 0.44 (+0.08) 3.39 (+3.03) ϵsharpsubscriptitalic-$\epsilon_{\text{sharp}}\downarrow$ 47.70 45.68 (-2.02) 46.82 (-0.88) 40.76 (-6.94) 44.13 (-3.57) 211.67 (+163.97)
Training Details. We train the ResNet18 model as described in Section C.1.2. We train using stochastic gradient descent (SGD) initialized with a 0.1 learning rate, 0.0005 weight decay, and 0.9 momentum. The learning rate then follows a cosine decay schedule, beginning at 0.1 with the same 0.0005 weight-decay setting. We have deployed the same configuration for measuring flatness as described in Section C.1.1 except for tinyImageNet experiment deploying $\tau=0.5,\epsilon=10.0$ for numerical stability.
C.1.5 Clean Accuracies of Models (Table 3)
    Dataset      ERM         AugMix         RandAug         PixMix         DeepAug         StyleAug ImageNet $\uparrow$ 69.48 69.33 (-0.15) 68.85 (-0.63) 66.86 (-2.62) 65.05 (-4.43) 54.35 (-15.13) tinyImageNet $\uparrow$ 64.74 65.08 (+0.34) 70.09 (+5.35) 66.97 (+2.23) 69.50 (+4.76) 50.60 (-14.14) CIFAR-10 $\uparrow$ 90.25 95.60 (+5.35) 95.77 (+5.52) 95.75 (+5.50) 80.92 (-9.33) 90.59 (+0.34) CIFAR-100 $\uparrow$ 64.82 70.40 (+5.58) 74.20 (+9.38) 74.68 (+9.86) 49.25 (-15.57) 38.76 (-26.06)
The tendency of the ’good’ and the ’bad’ augmentations remains the same as in Table 3, which aligns with our theorems, considering that generalization bound also applies to train-test data discrepancies (target data $\mathcal{T}$ being the test dataset.) On the other hand, there is a weak correlation between robustness and the clean accuracy. For instance, all methods have inferior clean accuracy compared to ERM in ImageNet, irrespective of their robustness. There are some cases where the clean accuracy increases but the robustness drops (e.g., StyleAug in CIFAR10), and some cases have dropped clean accuracy but improved robustness (e.g., PixMix in ImageNet). Also, some have severely dropped clean accuracy and robustness (e.g., StyleAug in CIFAR-100). We observe all four cases (better clean accuracy, better robustness / better clean accuracy, worse robustness / worse clean accuracy, better robustness / worse clean acc, worse robustness) exist.
C.1.6 Additional Experiments on Varying Backbone Networks
To validate our theoretical findings under differing backbone networks, we present additional experimental results on CIFAR-100 robustness benchmarks, extending results from Table 3. In CIFAR-100 benchmarks, AugMix, RandAug, and PixMix had nontrivial augmented sampled within the proximal region, leading to flatter minima and overall enhanced robustness. Contrarily, DeepAug and StyleAug had no representations near the original, failing to promote flatter minima and improved robustness. The tendency holds exactly the same for varying backbone networks as well. In below (Table  5), we demonstrate that our findings are consistent across different backbone networks [47, 23, 52].
Table 5: Comparison of various augmentation methods against ERM across different backbone networks on CIFAR-100 robustness benchmarks. ($\downarrow$: The lower, the better.)
Backbone Benchmark ERM      AugMix      RandAug      PixMix      DeepAug      StyleAug AllConvNet CIFAR-100-C $\downarrow$ 56.88 43.13 (-13.75) 47.52 (-9.36) 37.70 (-19.18) 62.00 (+5.12) 56.90 (+0.02) CIFAR-100-$\overline{\text{C}}$downarrow 20.33 9.84 (-10.49) 13.86 (-6.47) 14.29 (-6.04) 64.24 (+43.91) 38.64 (+18.31) CIFAR-100, $L_{2}$downarrow 93.96 86.51 (-7.45) 88.96 (-5.00) 74.93 (-19.03) 95.29 (+1.33) 96.56 (+2.60) CIFAR-100, $L_{\infty}$downarrow 99.95 99.90 (-0.05) 99.94 (-0.01) 99.64 (-0.31) 99.99 (+0.04) 99.71 (-0.24) DenseNet CIFAR-100-C $\downarrow$ 61.21 43.32 (-17.89) 48.54 (-12.67) 37.58 (-23.63) 61.47 (+0.26) 68.97 (+7.76) CIFAR-100-$\overline{\text{C}}$downarrow 23.10 8.87 (-14.23) 13.68 (-9.42) 14.27 (-8.83) 66.20 (+43.10) 43.90 (+20.80) CIFAR-100, $L_{2}$downarrow 99.91 97.88 (-2.03) 99.44 (-0.47) 93.19 (-6.72) 98.70 (-1.21) 99.76 (-0.15) CIFAR-100, $L_{\infty}$downarrow 100.00 99.99 (-0.01) 99.99 (-0.01) 99.99 (-0.01) 99.99 (-0.01) 99.99 (-0.01) ResNext29 CIFAR-100-C $\downarrow$ 58.05 39.60 (-18.45) 43.49 (-14.56) 31.35 (-26.70) 61.50 (+3.45) 70.11 (+12.06) CIFAR-100-$\overline{\text{C}}$downarrow 13.09 3.85 (-9.24) 6.99 (-6.10) 2.64 (-10.45) 66.61 (+53.52) 22.78 (+9.69) CIFAR-100, $L_{2}$downarrow 98.72 91.68 (-7.04) 95.95 (-2.77) 91.64 (-7.08) 98.36 (-0.36) 99.67 (+0.95) CIFAR-100, $L_{\infty}$downarrow 100.00 99.94 (-0.06) 99.97 (-0.03) 99.90 (-0.10) 99.96 (-0.04) 99.98 (-0.02)
Appendix D Detailed Review of Related Works
D.1 Data Augmentations Strategies
Existing augmentation methods can be grouped into model-free, model-based, and policy-based algorithms, borrowing the taxonomy suggested by [53].
Model-free augmentations can be divided into single-image and multiple-image methods. Single-image augmentations not only include elementary image manipulations such as translation, rotation, and color jittering but also includes masking methods such as CutOut [11] and Hide-and-Seek [31]. Multi-image augmentations exploit two or more images to construct augmented images, including the mixing of the two disjoint images and labels, such as Mixup [59] and CutMix [57]. Recently, it is empirically observed that a few methods, AugMix [21] and PixMix [22], exhibit their superiority over others in terms of model robustness.
Model-based augmentations utilize pretrained models to generate augmented data. A number of methods that exploit generative models including CGAN [40], and their variants [12, 38, 2, 54] aim to aid data imbalance problems. Other augmentations, including DeepAugment [19], ANT [44], and StyleAug [24], directly aim to enhance the classifier’s robustness against common corruption, adversarial attacks, or domain shifts.
Policy-based augmentations focus on designing an automatic way to determine the optimal augmentation strategies by employing reinforcement learning or adversarial training. From a pioneering method AutoAugment [9] that utilizes reinforcement learning for finding the best augmentation strategies, subsequent works such as Fast AA [36], Faster AA [18], and RandAugment [10] aim to enhance both the efficiency of policy search and the model performance. Adversarial training-based augmentation strategies, including AdaTransform [49], Adversarial AA [60], and AugMax [51] leverage adversarial perturbations which maximally disturbs samples to be misclassified into other labels, finally leading to improve model robustness against unseen domains.
Despite the effectiveness of augmentation methods in enhancing model performance, prior studies have focused more on their practical use rather than on understanding their theoretical impact on model robustness to data shifts.
D.2 Augmentations and Model Robustness
A number of previous works have tried to reveal the relationship between augmentation and model robustness.
[37] have provided a game-theoretic perspective on augmentations and introduces a new proxy metric for measuring common corruption robustness. [63] and [50] have theoretically found out that adversarial perturbations in the latent space can simulate worst-case distributional shifts in the data. [43] have empirically found out that Mixup and CutOut with model weight averaging is shown to improve adversarial robustness. [55] has shown that Mixup enlarges boundary thickness, the marginal space between differently labeled samples, which is deemed to be highly correlated with adversarial robustness and common corruptions. [56] interpret the augmentation-originated gains in model robustness by explaining that augmentations make deep models utilize both high and low frequency information of images so as to enhance model robustness against data corruptions. [41] and [1] have theoretically shown that utilizing unlabeled data in training can improve adversarial robustness. [19] have empirically found out that exploiting diverse augmentations together improves robustness against both adversarial and common corruptions.
Nonetheless, most of the prior interpretations of how augmentation contributes to model robustness are either confined to specific types of augmentations and robustness, or empirical analysis. None of the prior works generally explain how augmentations theoretically can improve model robustness across diverse distributional shifts.
D.3 Other Related Works
[61] theoretically show the duality between SAM and AT, and empirically found the conditions where SAM outperforms AT. Nonetheless, the study linked input–parameter duality only in binary linear classifiers with Gaussian data and basic adversarial-training (AT). On the other hand, we show that the duality survives in deep, multi-class models, arbitrary data distributions, and general augmentation family whose distribution is dense around the data. The findings of [61] cannot be extended to augmentation-driven robustness.
[27] investigates how to preserve fairness under distribution shifts by reinterpreting such shifts as equivalent perturbations to both model weights and inputs, employing a transportation function. However, their work focuses exclusively on group fairness with respect to demographic parity and equal opportunity, and only demonstrates the feasibility of this equivalence without further extensions.
[64] shows in binary classification that flatter minima mitigate the boundary tilting problem [48], and empirically finds these minima correlate with greater robustness to common corruptions. However, the theory remains limited, as the boundary tilting problem is closely tied to adversarial robustness and is not explored beyond binary classification.
Appendix E Further Discussions
E.1 Practical Applications of Our Theorems
Although the theoretical insights we provide into how data augmentation influences robustness are valuable in their own right, we herein present some practical consequences of our findings.
First, when proposing a novel, complex augmentation, prioritize densely populating the proximal region around the original image. In practice, researchers can perform a quick, low-cost check by inspecting how the augmented samples cluster near the source image. If that neighborhood is well covered, the resulting model is highly likely to remain robust under unforeseen data shifts.
Second, we can roughly forecast the benefit of combining two augmentations by looking at each augmentation’s sample density close to the original data. For example, DeepAug and StyleAug both leave the region empty, whereas AugMix fills it densely. Hence, pairing AugMix with either DeepAug or StyleAug should yield a substantial improvement (by filling the gap), while combining DeepAug with StyleAug alone should offer little extra robustness.
Figure 4: A 2D sample densities generated by each augmentation and their pairwise combinations. AugMix densely covers the proximal neighborhood, DeepAug and StyleAug leave it sparse, and mixing AugMix with either closes the gap. However, DeepAug + StyleAug still leaves a central hole.
Table 6: The effect of combining augmentations that have significant differences in proximal density. In all the cases, combining DeepAug or StyleAug with AugMix yielded notable robustness gain. Nevertheless, combining DeepAug with StyleAug yielded marginal improvement.
Benchmark AugMix DeepAug StyleAug DeepAug + AugMix StyleAug + AugMix DeepAug + StyleAug CIFAR-10-C $\downarrow$ 15.24 31.92 30.50 15.04 (-16.88) 19.02 (-11.48) 24.84 (-5.66) CIFAR-10-$\overline{\text{C}}$downarrow 1.73 17.99 9.34 1.32 (-16.67) 1.55 (-7.79) 4.12 (-5.22) CIFAR-100-C $\downarrow$ 42.64 62.34 70.92 43.53 (-18.81) 49.13 (-21.79) 53.50 (-17.42) CIFAR-100-$\overline{\text{C}}$downarrow 5.01 39.54 30.77 3.67 (-35.87) 3.70 (-27.07) 7.60 (-23.17) CIFAR-10, $L_{2}$downarrow 70.68 91.07 91.13 69.30 (-21.77) 81.07 (-10.06) 87.30 (-3.83) CIFAR-10, $L_{\infty}$downarrow 99.57 100.00 99.97 94.15 (-4.22) 96.34 (-3.33) 99.80 (-0.17) CIFAR-100, $L_{2}$downarrow 92.78 98.37 99.67 94.15 (-4.22) 96.34 (-3.33) 97.53 (-2.14) CIFAR-100, $L_{\infty}$downarrow 99.67 99.99 99.98 99.89 (-0.10) 99.95 (-0.03) 99.96 (-0.02)
Our study also provides explanations on the behavior of adversarial training (AT), and rooms for the improvement. While AT is known to strengthen robustness to adversarial attacks, models trained with AT still break down when the test distribution deviates largely from the training data, as in common corruption case [20, 39]. We conjecture that this fragility arises because AT typically discovers small $b$-flat minima: these minima are flat enough to absorb minor, norm-bounded perturbations, yet they fail to cope with large distribution shifts. Hence, pairing AT with data augmentation schemes that generate samples both close to and far from the original data manifold (such as AugMix) can offset AT’s weakness under severe shifts. Below (Table 88), we report results obtained by training with AT (FGSM, $\epsilon=8/255$) and AT in conjunction with AugMix.
Table 7: Flatness evaluations of AT and AT + AugMix on CIFAR-10/100.
Dataset Metrics AT AT + AugMix CIFAR-10 $\mu_{\text{PAC-Bayes}}\downarrow$ 158.15 113.78 (-44.37) LPF $\downarrow$ 1.25 0.58 (-0.67) $\epsilon_{\text{sharp}}$downarrow 16.36 10.29 (-6.07) CIFAR-100 $\mu_{\text{PAC-Bayes}}\downarrow$ 207.90 161.89 (-46.01) LPF $\downarrow$ 2.95 2.11 (-0.84) ϵsharpsubscriptitalic-$\epsilon_{\text{sharp}}\downarrow$ 20.81 10.34 (-10.47)
Table 8: CIFAR-10/100 robustness benchmark results of AT and AT + AugMix.
Robustness Benchmark AT AT + AugMix CIFAR-10-C $\downarrow$ 24.12 15.39 (-8.73) Common CIFAR-10-$\overline{\text{C}}$downarrow 3.25 1.30 (-1.95) Corruption CIFAR-100-C $\downarrow$ 52.50 44.03 (-8.47) CIFAR-100-$\overline{\text{C}}$downarrow 7.36 3.74 (-3.62) CIFAR-10, $L_{2}$downarrow 69.24 62.86 (-6.38) Adversarial CIFAR-10, $L_{\infty}$downarrow 94.55 98.40 (+3.85) Attacks CIFAR-100, $L_{2}$downarrow 89.28 88.60 (-0.68) CIFAR-100, $L_{\infty}$downarrow 98.42 99.75 (+1.33)
E.2 Comparison with Flat Minima Optimizers
We establish a formal connection between data augmentation and robustness through the notion of flat minima. Specifically, we prove the duality between the input space and parameter space manifolds (Theorem 1) and show that the nonzero probability proximal region induced by augmentation (Assumption 1) is mapped to an equivalent region in parameter space, promoting flatter minima (Theorem 2). Bridging these augmentation-induced flat minima to improved robustness (Theorem 3) leverages prior results indicating that flat minima tighten generalization bounds [7, 62] against distribution shifts. Empirically, several studies already confirm that optimizers explicitly seeking flat minima enhance robustness [15, 5, 28, 33, 61]. Nevertheless, our findings are not trivial extensions of the prior works on flat optimizers. The experiments reveal that not every augmentation leads to flatter minima or greater robustness. Our work provides a necessary condition for an augmentation to foster flatter minima (non-zero proximal density). This property correlates strongly with flat minima and robustness in our experiments. Together with the proven duality between input and parameter manifolds under general theoretical setups, we provide a distinct and novel contribution apart from existing research on flat minima optimizations.
E.3 Any Other Modalities Beyond Visual Data?
Our main experiments handle visual data, more specifically, image classification tasks. Also, it might be unexplored that the key claims of this work can be further applied to other modalities, such as language, audio, or even sensory signals from devices. Our experiments seem to be limited to vision cases, but we believe that it is an important step forward to understand theoretical importance of augmentations in machine learning. Furthermore, we want to emphasize that our main theoretical claims do not explicitly rely on the characteristics of visual data, so it can be further extended to the other form of modalities. In addition, extending to other complicated task beyond classification can be an appropriate future research direction. In particular, sequential training such as natural language processing or decision based on reinforcement learning probably show distinct and unique theoretical understanding beyond classification.